OpenStudy (anonymous):

using the fact that sin47=0.7314, evaluate cos223

OpenStudy (anonymous):

i get -0.7314 is this correct?

OpenStudy (raden):

cos(270 - x) = -sin(x) so, cos(270 - 47) = -sin47 = -0.7314, you are right !

OpenStudy (anonymous):

thank you :)

OpenStudy (raden):

welcome :)